OpenStudy (lena772):

The radius of a cone-shaped tank is 4 feet less than its height. If the height of the tank is (x - 3) feet, the expression below shows the volume of the tank. 1 over 3π (x - 7)2 (x - 3) What does the factor π(x - 7)2 (x - 3) represent? The area of the base of the tank The area of the curved sides of the tank The volume of 3 of the same cone-shaped tanks The volume of 6 of the same cone-shaped tanks

OpenStudy (lena772):

@AllTehMaffs

OpenStudy (anonymous):

hi hi - what's your thought? ^^

OpenStudy (lena772):

I have no idea

OpenStudy (lena772):

My guess would be A

OpenStudy (anonymous):

if \[V= \frac{1}{3}\pi(x-7)^2(x-3) =\ \text{Volume of one cone}\] then \[ \pi(x-7)^2(x-3) = 3 \Big(\frac{1}{3}\pi(x-7)^2(x-3) \Big) = 3V= \ \text{Volume of ? cones}\]

OpenStudy (lena772):

3

OpenStudy (anonymous):

yah ^_^

OpenStudy (anonymous):

So the answer would be C! Do you know where those expressions they used came from?

OpenStudy (lena772):

No

OpenStudy (anonymous):

The volume of a cone is \[V_{cone}= \frac{1}{3} BH\] Its base is the area of a circle \[ B=\pi R^2\] and its height is just H. So in this problem, the height was given by \[H=(x-3)ft\] and the radius of the base was "4 feet less than the height" \[R=H-4ft = (x-3)ft-4ft = (x-7) ft\] So the volume expression that they gave would be \[V_{cone}= \frac{1}{3} (B)(H) =\frac{1}{3} (\pi R^2)(H) = \frac{1}{3} \Big(\pi(x-7)^2\Big)\Big(x-3\Big)ft^3\]